# Recursive Quantum Relaxation for Combinatorial Optimization Problems

### v2

@@ Abstract
Quantum optimization methods use a continuous degree-of-freedom of quantum states to heuristically solve combinatorial problems, such as the MAX-CUT problem, which can be attributed to various NP-hard combinatorial problems. This paper shows that some existing quantum optimization methods can be unified into a solver that finds the binary solution that is most likely measured from the optimal quantum state. Combining this finding with the concept of quantum random access codes (QRACs) for encoding bits into quantum states on fewer qubits, we propose an efficient recursive quantum relaxation method called recursive quantum random access optimization (RQRAO) for MAX-CUT. Experiments on standard benchmark graphs with several hundred nodes in the MAX-CUT problem, conducted in a fully classical manner using a tensor network technique, show that RQRAO outperforms the Goemans–Williamson method and is comparable to state-of-the-art classical solvers. The codes will be made available soon.
@@ 1 Introduction
Figure 1: Concept of the proposed method, recursive quantum random access optimization (RQRAO). (a) One/two/three classical bits are embedded into one qubit using a quantum random access code (QRAC) [1, 2]. A total of two/four/eight variants of the information of the one/two/three classical bits are represented as a vector pointing to the corner of the line/square/cube inscribed in the Bloch sphere. These specific quantum states are called (one-qubit) magic states. (b) Using a high compression QRAC, the size of the Hilbert space corresponding to the given problem is reduced and each (multi-qubit) magic state is no longer necessarily orthogonal to the others. Note that only four of the $2^{6}$ magic states are illustrated. After optimizing the objective function defined on the reduced Hilbert space, the bit string is decoded by searching for the nearest magic state in the fidelity, which is called rounding. To accomplish the rounding without strong approximation, a recursive strategy similar to the RQAOA [3] is adopted.
@@
The MAX-CUT problem is one of Karp’s 21 NP-complete problems [4] and is frequently used as a benchmark of combinatorial optimization algorithms because of its simplicity. Namely, given a graph that consists of nodes connected by weighted edges, the optimization variant of the MAX-CUT problem asks for the labeling of nodes into two classes such that the sum of the weighted edges connecting nodes with different labels is maximized. Obtaining good solutions for the MAX-CUT problem is in great demand, because various NP-hard combinatorial problems, e.g., partitioning, covering, and coloring problems, can be relaxed to the MAX-CUT problem by treating the constraints as penalty terms [5, 6]. Although the MAX-CUT problem is a discrete optimization problem, it can be translated to a continuous variant by treating the discrete solution in a probabilistic manner. Classically, such a continuous variant of the MAX-CUT problem is relaxed to a semidefinite program by approximating the probability [7]. However, the quantum optimization does not take such an approximation because the probabilistic characteristics are naturally incorporated into the quantum states. In the quantum MAX-CUT , all solutions are encoded in the eigenstates of the problem Hamiltonian, and then the quantum state is optimized such that the expected value of the problem Hamiltonian is maximized. From a theoretical viewpoint, it has been proved that the eigenstate with the maximum eigenvalue can be obtained by the infinite-level quantum approximate optimization algorithm (QAOA) [8, 9] with the appropriate parameters. However, in a noisy quantum device, it is intractable to carry out even a sufficiently large level QAOA, let alone an infinite level QAOA. For this reason, the field of combinatorial problems with quantum computing is dominated by the development of heuristics, that mainly focus on the development of the ansatz variants and improving the trainability of QAOA [10].
@@
Rather than developing the QAOA ansatz, an alternative approach is to develop a non-diagonal Hamiltonian instead of the diagonal Ising-type one used in the QAOA. In [11], the MAX-CUT Hamiltonian is constructed using quantum random access codes (QRACs) [1, 2] (see Sec. 5.3 for a review). Using QRACs, the $n$-qubit quantum state can be used to encode $m$ binary variables for $m>n$. The QRAC Hamiltonian is still a 2-local Hamiltonian, and therefore it is QMA-hard to find the ground state [12], but the QRAC formulation has the advantage that the number of qubits can be reduced, making it easier to optimize the variational state.
@@
    We theoretically show that incorporating QRACs into the RQAOA framework is a natural expansion of [11] and [3], i.e., they are all uniformly expressed as a method that (i) maximizes the expectation of the problem Hamiltonian that all candidate solutions are embedded in using QRACs, and (ii) searches for the nearest embedding in the fidelity (Theorem 1).
@@
The remainder of the paper is organized as follows: Section 2 presents the mathematical definition of the MAX-CUT problem and describes the RQAOA and QRAC in brief. Section 3 describes the proposed algorithm and its theoretical background. Section 4 presents the results of numerical experiments. Section 5 summarizes the related works. Section 6 concludes the paper.
@@ 2.1 MAX-CUT Problem
is the cut weight corresponding to ${\bf b}$.
@@ 2.2 Recursive QAOA
Recursive QAOA (RQAOA) [3] is a variant of QAOA, which is an iterative method that determines the parity of one edge per iteration, then deletes a node and modifies a graph. Here, edge parity is defined as follows:
@@
Given edge $(j,k)$, positive and negative edges are defined as $b_{j}=b_{k}$ and $b_{j}\neq b_{k}$, respectively, where $b_{i}\in\{0,1\}$ is an attribute of node $i$. Define edge parity as being a positive or negative edge.
@@
The number of nodes is gradually decreased by repeating the edge parity determination and graph size reduction. When the number of nodes becomes smaller than a threshold value $M$, the parities of the remaining nodes are determined by brute-force search.
@@ 2.3 Quantum Random Access Optimization
In the Quantum Random Access Optimization (QRAO) [11], binary node attributes are encoded in fewer qubits than the number of nodes using quantum random cccess code (QRAC) [1, 2]. For each node $j\in V$, binary node attribute $b_{j}\in\{0,1\}$ is randomly mapped to $(q^{(j)},P^{(j)})$, where $q^{(j)}$ is a qubit index and $P^{(j)}$ is a Pauli matrix with the following constraints:
@@
where $P_{\langle j\rangle}:=P^{(j)}_{q^{(j)}}$ is a 1-local Pauli matrix corresponding to the $q^{(j)}$th qubit. If the given graph is sparse, constraint Eq. (9)${}_{2}$ rarely restricts the assignment, and hence, the total number of qubits can be reduced by up to $1/m$th the number of nodes.
@@
After maximizing ${\rm tr}(H_{m}\rho)$ with respect to $\rho$, the classical bit string ${\bf b}$ is decoded from the optimized $\rho$. This process is called rounding. In [11], two kinds of rounding were proposed, magic state rounding and Pauli rounding. Note that magic state rounding is a measurement method rather than a decoding method, and hence magic state rounding is henceforth referred to as random magic measurements. The random magic measurement is a projective measurement whose basis is the magic state selected uniformly at random. The detailed protocol is defined in Definition 3 of Supplementary Information A. Note that $(1,1)$-random magic measurement with $\mathfrak{P}=\{Z\}$ is equivalent to measurements on a computational basis. Pauli rounding is a decoding method based on the value of ${\rm tr}(P_{\langle j\rangle}\rho)$. The value of ${\bf b}$ is determined as $b_{j}=0$ if ${\rm tr}(P_{\langle j\rangle}\rho)>0$ and $b_{j}=1$ if ${\rm tr}(P_{\langle j\rangle}\rho)<0$. If ${\rm tr}(P_{\langle j\rangle}\rho)=0$, $b_{j}$ is determined as 00 or $1$ uniformly at random. In Sec. 3.2, we provide theorems to show the relationship between random magic measurements and Pauli rounding and reveal what assumption is made in Pauli rounding.
@@ 3 Proposed Method
Section 3.1 describes the concrete algorithm of the proposed method. In brief, the proposed method incorporates the $(m,1)$-QRAC Hamiltonian (Eq. (10)) into the RQAOA framework (Sec. 2.2). The edge parity is recursively determined based on the value of the edge energy, which is defined using the $(m,1)$-QRAC Hamiltonian. Two heuristics are also incorporated into the proposed method. One is the ensemble method of defining the edge energy, and the other determines several edge parities at once using the maximum spanning tree. The former improves the resulting cut weight whereas the latter reduces the runtime of the proposed method.
In Sec. 3.2, we show that not only the proposed method but also RQAOA and QRAO
@@
    try to maximize the expected cut weight using $(m,1)$-random magic measurements with respect to the quantum state, and
@@
    try to decode the bit string embedded in the nearest magic state in the view of the fidelity against the optimized quantum state.
@@ 19             $G^{\prime}=(V^{\prime},E^{\prime})\leftarrow\texttt{GetReducedGraph}(G^{\prime},(j,k))$
Add small perturbation noise to all edge weights to mitigate the isolated nodes problem (see Supplementary Information B).
@@
Line 5: Optimization
Define ansatz $|\psi({\boldsymbol{\theta}})\rangle$, where ${\boldsymbol{\theta}}\in\mathbb{R}^{K}$ are $K$ learnable parameters. The objective function is defined as
@@
The optimized quantum state $|\psi({\boldsymbol{\theta}}^{*})\rangle$ is obtained by maximizing $\mathcal{L}^{(t)}({\boldsymbol{\theta}})$. Any optimization method can be used here.
@@
Lines 7–11: Compute the Ensemble Edge Energy
@@
where $\mu_{jk}$ and $\sigma_{jk}$ are the mean and standard deviation of $\{\mathcal{E}^{(t)}_{jk}\}^{N}_{t=1}$, which are defined as $\mu_{jk}:=\frac{1}{N}\sum^{N}_{t=1}\mathcal{E}^{(t)}_{jk}$ and $\sigma_{jk}:=\sqrt{\frac{1}{N}\sum^{N}_{t=1}(\mathcal{E}^{(t)}_{jk}-\mu_{jk})^{2}}$, respectively. Equation (17) means that $\mathfrak{E}_{jk}$ has a nonzero value if $\mu_{jk}\pm S\sigma_{jk}$ have the same sign as $\mu_{jk}$.
Lines 12–16: Compute the Maximum Spanning Tree
@@
Lines 17–21: Modify the Graph
@@
Line 23: Brute-force Search
The recursive procedure is stopped when the number of nodes of $G^{\prime}$ is less than or equal to $M$. Here, the number of nodes for which parity with other nodes has not yet been determined is less than or equal to $M$. That is, the number of candidate solutions is less than or equal to $2^{M}$. If $M$ is sufficiently small, e.g., $M\leq 20$, a brute-force search among the candidates is not time-consuming, and hence the best solution ${\bf b}^{*}$ can be quickly obtained. The RQAOA [3] also uses brute-force search.
@@
The whole RQRAO procedure is schematically shown in Fig. 4. The problem graph $G$ is shown in Fig. 4, where the red and blue edges have weights $+1$ and $-1$, respectively. The random Pauli assignments and optimization are carried out to obtain the edge energy for each trial $t$, $\mathcal{E}^{(t)}_{jk}$ (Fig. 4. The ensemble edge energy $\mathfrak{E}_{jk}$ is then calculated using $\{\mathcal{E}^{(t)}_{jk}\}^{N}_{t=1}$ (Fig. 4. The inconsistent edges are removed by solving the maximum spanning tree problem, where $|\mathfrak{E}_{jk}|$ are regarded as the edge weights. The rooted tree is constructed by giving the direction from root to leaf, where the root node has been randomly selected (Fig. 4. The edge parities in the rooted tree are determined from leaf to root according to the sign of corresponding $\mathfrak{E}_{jk}$. At the same time, the graph is modified to satisfy the determined parity. Figure 4 represents the modified graph after all edge parities in the rooted tree have been determined. The protocol from Fig. 4 to 4 is repeated until the number of nodes of the modified graph becomes smaller than $M$.

### v3

@@ Abstract
Quantum optimization methods use a continuous degree-of-freedom of quantum states to heuristically solve combinatorial problems, such as the MAX-CUT problem, which can be attributed to various NP-hard combinatorial problems. This paper shows that some existing quantum optimization methods can be unified into a solver to find the binary solution which is most likely measured from the optimal quantum state. Combining this finding with the concept of quantum random access codes (QRACs) for encoding bits into quantum states on fewer qubits, we propose an efficient recursive quantum relaxation method called recursive quantum random access optimization (RQRAO) for MAX-CUT. Experiments on standard benchmark graphs with several hundred nodes in the MAX-CUT problem, conducted in a fully classical manner using a tensor network technique, show that RQRAO not only outperforms the Goemans–Williamson and recursive QAOA methods, but also is comparable to state-of-the-art classical solvers. The code is available at https://github.com/ToyotaCRDL/rqrao.
@@ 1 Introduction
Figure 1: Concept of the proposed method, recursive quantum random access optimization (RQRAO). (a) One/two/three classical bits are embedded into one qubit using a quantum random access code (QRAC) [1, 2]. A total of two/four/eight variants of the information of the one/two/three classical bits are represented as a vector pointing to the corner of the line/square/cube inscribed in the Bloch sphere. These specific quantum states are called (one-qubit) magic states. (b) Using a high compression QRAC, the size of the Hilbert space corresponding to the given problem is reduced and each (multi-qubit) magic state is no longer necessarily orthogonal to the others. Note that only four of the $2^{6}$ magic states are illustrated. After optimizing the objective function defined on the reduced Hilbert space, the bit string is decoded by searching for the nearest magic state as measured in the fidelity, which is called rounding. To accomplish a good rounding, a recursive strategy similar to the RQAOA [3] is adopted.
@@
The MAX-CUT problem is one of Karp’s 21 NP-complete problems [4] and is frequently used as a benchmark of combinatorial optimization algorithms because of its simplicity. Namely, given a graph that consists of nodes connected by weighted edges, the optimization variant of the MAX-CUT problem asks for the labeling of nodes into two classes such that the sum of the weighted edges connecting nodes with different labels is maximized. Obtaining good solutions for the MAX-CUT problem is in great demand, because various NP-hard combinatorial problems, e.g., partitioning, covering, and coloring problems, can be transformed into the MAX-CUT problem [5, 6]. Although the MAX-CUT problem is a discrete optimization problem, it can be solved by relaxing the discrete variables into continuous ones. For example, the semidefinite program in [7] finds an embedding of binary variables as real-valued vectors in a high-dimensional sphere so that the labels can be probabilistically determined by partitioning the vectors with a random hyperplane cutting. (see Supplementary Information H.1). However, the quantum optimization does not take such an approximation because the probabilistic characteristics are naturally incorporated into the quantum states. In most approaches to the MAX-CUT problem using quantum algorithms, all solutions are encoded in the eigenstates of the problem Hamiltonian, and then the quantum state is optimized such that the expected value of the problem Hamiltonian is maximized (see Supplementary Information H.2). From a theoretical viewpoint, it has been proven that the eigenstate with the maximum eigenvalue can be obtained by the infinite-level quantum approximate optimization algorithm (QAOA) [8, 9] with the appropriate parameters. However, in a noisy quantum device, it is intractable to carry out even a sufficiently large level QAOA, let alone an infinite level QAOA. For this reason, the field of combinatorial problems with quantum computing is dominated by the development of heuristics, that mainly focus on the development of the ansatz variants and improving the trainability of QAOA [10].
@@
Rather than developing the QAOA ansatz, an alternative approach is to develop a non-diagonal Hamiltonian instead of the diagonal Ising-type one used in the QAOA. In [11], the MAX-CUT Hamiltonian is constructed using quantum random access codes (QRACs) [1, 2] (see Sec. 2.3 for details, and Supplementary Information H.3 for a review). Using QRACs, the $n$-qubit quantum state can be used to encode $m$ binary variables for $m>n$. The QRAC Hamiltonian is still a 2-local Hamiltonian, and therefore it is QMA-hard to find the ground state [12]. However, the QRAC formulation has the advantage that the number of qubits can be reduced, making it easier to optimize the variational state.
@@
    We theoretically show that incorporating QRACs into the RQAOA framework is a natural extension of [11] and [3], i.e., they are all uniformly expressed as a method that (i) maximizes the expectation of the problem Hamiltonian so that all candidate solutions are embedded using QRACs, and (ii) searches for the nearest embedding as measured in the fidelity (Theorem 1).
@@
Note that the proposed framework is entirely classical except for the ground state preparation step, and therefore it may not provide the so-called quantum advantage. However, when an efficient quantum algorithm is used for the ground-state preparation of the framework, the overall runtime can be improved.
The remainder of the paper is organized as follows: Section 2 presents the mathematical definition of the MAX-CUT problem and describes the RQAOA and QRAC in brief. Section 3 describes the proposed algorithm and its theoretical background. Section 4 presents the results of numerical experiments. Section 5 concludes the paper.
@@ 2.1 MAX-CUT Problem
is the cut weight corresponding to ${\bf b}=(b_{0},\ldots,b_{|V|-1})$, where $b_{i}\in\{0,1\}$ is an attribute of node $i$.
@@ 2.2 Recursive QAOA
Recursive QAOA (RQAOA) [3] is a variant of QAOA for the MAX-CUT problem, which is an iterative method that determines the parity of one edge in a graph per iteration, deletes a node connected by the edge and modifies the graph. Here, edge parity is defined as follows:
@@
Given edge $(j,k)$, positive and negative edges are defined as $b_{j}=b_{k}$ and $b_{j}\neq b_{k}$, respectively. The edge parity is said positive (negative) for a positive (negative) edge.
@@
The number of nodes is gradually decreased by repeating the edge parity determination and graph size reduction. When the number of nodes becomes smaller than a threshold value $M$, the parities of the remaining nodes are determined by brute-force search. Once all parities are determined, a label for one node is randomly set to either 0 or 1. Subsequently, the labels for all other nodes are determined based on the assigned parities.
@@ 2.3 Quantum Random Access Optimization
In the Quantum Random Access Optimization (QRAO) [11], binary node attributes are encoded in fewer qubits than the number of nodes using quantum random access code (QRAC) [1, 2]. For each node $j\in V$, binary node attribute $b_{j}\in\{0,1\}$ is randomly mapped to $(q^{(j)},P^{(j)})$, where $q^{(j)}$ is a qubit index and $P^{(j)}$ is a Pauli matrix with the following constraints:
@@
where $P_{\langle j\rangle}:=P^{(j)}_{q^{(j)}}$ is a 1-local Pauli matrix corresponding to the $q^{(j)}$th qubit. If the given graph is sparse, the second constraint in Eq. (9) rarely restricts the assignment, and hence, the total number of qubits can be reduced by up to $1/m$-th the number of nodes.
@@
After maximizing ${\rm tr}(H_{m}\rho)$ with respect to $\rho$, the classical bit string ${\bf b}$ is decoded from the optimized $\rho$. This process is called rounding. In [11], two kinds of rounding were proposed, magic state rounding and Pauli rounding. Note that magic state rounding is the projective measurement whose basis is the magic state selected uniformly at random. The detailed protocol is defined in Definition 3 of Supplementary Information A, and as can be seen to closely resemble the classical shadow with Random Pauli measurements [17]. For this reason, the magic state rounding is henceforth referred to as random magic measurement. Note that $(1,1)$-random magic measurement with $\mathfrak{P}=\{Z\}$ is equivalent to measurements on a computational basis. Pauli rounding is a decoding method based on the value of ${\rm tr}(P_{\langle j\rangle}\rho)$. The value of ${\bf b}$ is determined as
$b_{j}=\left\{\begin{array}[]{ll}0&\big{(}{\rm tr}(P_{\langle j\rangle}\rho)>0\big{)}\\[5.69054pt] 1&\big{(}{\rm tr}(P_{\langle j\rangle}\rho)<0\big{)}\end{array}\right..$ (14)
If ${\rm tr}(P_{\langle j\rangle}\rho)=0$, $b_{j}$ is determined as 00 or $1$ uniformly at random. In Sec. 3.2, we provide theorems to show the relationship between random magic measurements and Pauli rounding and reveal what assumption is made in Pauli rounding.
@@ 3 Proposed Method
Section 3.1 describes the concrete algorithm of the proposed method. In brief, the proposed method incorporates the $(m,1)$-QRAC Hamiltonian (Eq. (10)) into the RQAOA framework (Sec. 2.2). The edge parity is recursively determined based on the value of the edge energy, which is defined using the $(m,1)$-QRAC Hamiltonian. Two new heuristics are also incorporated into the proposed method: the ensemble method of defining the edge energy, and determining several edge parities at once using the maximum spanning tree. The former improves the resulting cut weight whereas the latter reduces the runtime of the proposed method.
In Sec. 3.2, we show that the proposed method, RQAOA, and QRAO
@@
    maximize the expected cut weight using $(m,1)$-random magic measurements with respect to the quantum state, and
@@
    decode the bit string embedded in the nearest magic state as measured in the fidelity against the optimized quantum state.
@@ 19             $G^{\prime}=(V^{\prime},E^{\prime})\leftarrow\texttt{GetReducedGraph}(G^{\prime},(j,k))$
Add small perturbation noise to all edge weights to mitigate the isolated nodes problem (see Supplementary Information B). We note that the isolated nodes problem also occurs at RQAOA [3] although not reported in the previous literature. We found out the perturbation is effective to deal with the problem.
@@
Line 5: Optimize the Quantum State
Define ansatz $|\psi({\boldsymbol{\theta}})\rangle$, where ${\boldsymbol{\theta}}\in\mathbb{R}^{K}$ are $K$ learnable parameters. For example, we used the matrix product state ansatz in the numerical experiments, where all matrix elements were treated as the learnable parameters (see Implementation Details in Sec. 4). The objective function is defined as
@@
The optimized quantum state $|\psi({\boldsymbol{\theta}}^{*})\rangle$ is obtained by maximizing $\mathcal{L}^{(t)}({\boldsymbol{\theta}})$. Any optimization method for the parameters of the ansatz can be used here.
@@
Lines 8–12: Compute the Ensemble Edge Energy
@@
where $\mu_{jk}$ and $\sigma_{jk}$ are the mean and standard deviation of $\{\mathcal{E}^{(t)}_{jk}\}^{N}_{t=1}$, which are defined as $\mu_{jk}:=\frac{1}{N}\sum^{N}_{t=1}\mathcal{E}^{(t)}_{jk}$ and $\sigma_{jk}:=\sqrt{\frac{1}{N}\sum^{N}_{t=1}(\mathcal{E}^{(t)}_{jk}-\mu_{jk})^{2}}$, respectively , and $S$ is a hyperparameter designed such that if $|\mu_{jk}|\leq S\sigma_{jk}$, then $\mathfrak{E}_{jk}=0$.
The reason for taking an ensemble is twofold. First, unlike QAOA whose edge energy is solely from the ZZ interaction as in Eq. (4), the edge energy of QRAO can be defined by 9 different pairs of 1-local Pauli as in Eq. (17) whose corresponding Hamiltonian and ground state can be significantly different: some may be easier to compute with certain ansatz than others. The ensemble is utilized to take advantage of the various Hamiltonians and ground states. Second, taking an ensemble average of the edge energy from different pairs of 1-local Pauli is akin to bagging [18] in machine learning. Namely, even if an individual pair of 1-local Pauli is a weak predictor of the edge parity, their averages can give higher predictive performance when they are weakly correlated with each other. In our method, when optimization is performed on each of the $N$ different $(m,1)$-QRAC Hamiltonians created from different Pauli assignments, $N$ quantum states are obtained. Since these quantum states are optimized with respect to different Hamiltonians as objective functions, they can be considered analogous to the weak predictors mentioned above. Therefore, we expect an ensemble edge energy from these quantum states will yield a stronger predictor of the edge parity.
Line 14: Compute the Maximum Spanning Tree
@@
Lines 15–18: Modify the Graph
@@
Figure 4: Schematic illustration of tree construction and node deletion ordering: (a) Solid and dashed lines indicate edges with energy above and below the threshold, respectively; (b) Trees formed from the solid lines in (a); (c) Root nodes are randomly selected for each tree, and the nodes to be deleted are ordered from the leaf to the root.
Here, we note that the proposed method differs from the RQAOA in two parts: on which nodes and on how many nodes to be removed per iteration. In RQAOA, only one node at one end of the edge with the maximum absolute edge energy is removed per iteration, which is essentialy the only choice up to symmetricity of the ZZ interaction. In contrast, in the proposed method, any node that is an endpoint of an edge whose absolute edge energy exceeding the threshold is a candidate for removal per iteration (Fig. 4). To perform this multiple node removal without inconsistency, some edges that form a cycle are removed by calculating a spanning tree. If an edge is not part of the spanning tree, it is ignored, even if its edge energy exceeds the threshold. In the step where spanning trees are formed from the edges exceeding the threshold, a tree is created for each disconnected set of edges (Fig. 4). For each tree, a rooted tree is created by selecting a node at random as the root (Fig. 4), and node removal is performed from the leaves toward the root.
Line 20: Brute-force Search
This is essentialy the same as in RQAOA [3]: the recursive procedure is stopped when the number of nodes of $G^{\prime}$ is less than or equal to some constant $M$. In such case, the number of nodes whose parities are undetermined is at most $M$, and therefore the number of candidate solutions is at most $2^{M}$. The best solution ${\bf b}^{*}$ can be quickly obtained by a brute-force search.
@@
The whole RQRAO procedure is schematically shown in Fig. 5. The problem graph $G$ is shown in Fig. 5, where the red and blue edges have weights $+1$ and $-1$, respectively. The random Pauli assignments and optimization are carried out to obtain the edge energy for each trial $t$, $\mathcal{E}^{(t)}_{jk}$ (Fig. 5). The ensemble edge energy $\mathfrak{E}_{jk}$ is then calculated using $\{\mathcal{E}^{(t)}_{jk}\}^{N}_{t=1}$ (Fig. 5). The inconsistent edges are removed by solving the maximum spanning tree problem, where $|\mathfrak{E}_{jk}|$ are regarded as the edge weights. The rooted tree is constructed by giving the direction from root to leaf, where the root node has been randomly selected (Fig. 5). The edge parities in the rooted tree are determined from leaf to root according to the sign of corresponding $\mathfrak{E}_{jk}$. At the same time, the graph is modified to satisfy the determined parity. Figure 5 represents the modified graph after all edge parities in the rooted tree have been determined. The protocol from Fig. 5 to 5 is repeated until the number of nodes of the modified graph becomes at most $M$.